Incomplete manufacturing costs, expenses, and selling data for two different cases are as follows.

Incomplete manufacturing costs, expenses, and sell
Incomplete manufacturing costs, expenses, and sell

Collapse question part

(a)

Indicate the missing amount for each letter.

Case

1

2

Direct materials used $9,600

$

Direct labor 5,000 8,000
Manufacturing overhead 8,000 4,000
Total manufacturing costs

16,000
Beginning work in process inventory 1,000

Ending work in process inventory

3,000
Sales revenue 24,500

Sales discounts 2,500 1,400
Cost of goods manufactured 17,000 22,000
Beginning finished goods inventory

3,300
Goods available for sale 20,000

Cost of goods sold

Ending finished goods inventory 3,400 2,500
Gross profit

7,000
Operating expenses 2,500

Net income

5,000

b) Prepare a condensed cost of goods manaufactures schdule ffor Case1

c1) Prepare an Income statement for Case 1

c2) Prepare the current assets section of the balance sheet for Case1. Assume that in Case 1 the other items in the current assets section are as follows: Cash $4,000, Receivables (net) $15,000, Raw Materials $600 and Prepaid Expenses $400.
Answer #1
Concepts and reason

Cost: The estimated figure of various expenses related to the production of the merchandise is cost. The cost which has a direct relation to the product is a direct cost. The product which has an indirect relation to the product is an indirect cost. Both costs include the cost of material, labor, and expenses.

Fundamentals

Direct material: It is one of the product cost which has a direct relation to the manufacture of the product. The raw material is the basic requirement to produce the finished goods. The cost of direct material can be directly chargeable to the final product as compared to the indirect material.

Direct labor: It is one of the conversion cost which is required to transform the raw material inventory into finished goods inventory. Direct labor can be directly traceable to a specific product or cost center. In other words, the cost of the workers which is directly involved in the transformation of the raw material into the finished product.

Manufacturing overheads: It is an indirect cost which is required for production. It includes the cost of indirect material, indirect labor, and indirect expenses. The expenses such as depreciation on machines, foreman salary, electricity expenses of the factory are classified under manufacturing overhead.

Manufacturing cost: The cost and expenses incurred to produce the inventory are included in manufacturing cost. It includes the prime cost and manufacturing overheads. The prime cost is the first cost incurred on production and it includes the cost of direct material, cost of direct labor and cost of direct expenses.

Work-in-progress inventory: The inventory which has been finished in production are work-in-progress inventory. The process of production has applied to the raw material but it has not been completed. It includes the amount of direct material, direct labor, and expenses which are incurred in the production of this inventory till yet.

Sales: It is the main revenue of the organization. Various expenses, fixed expenses and adjustment of inventory to be done from the sales revenue. It is computed by multiplying the units sold and selling price per unit.

Finished goods inventory: The inventory which is completed in production and available for sale. The finished goods cost includes the prime cost, manufacturing overheads, and administration overheads.

Gross profit: It is the first profit earned after reducing the cost of inventory sold from the sales. The non-operating expenses are not yet reduced from the gross profit.

Operating expenses: The cost and expenses which are required to operate the business and generate revenue.

Net income: It is the income earned by reducing all operating and non-operating expenses from the sales revenue. Alternatively, non-operating expenses to be deducted from gross profit to determine net income.

(a.1)

Compute total manufacturing cost of case 1 as given below:

Totalmanufacturingcost=(Directmaterial)+(Directlabor)+(Manufacturingoverhead)=$9,600+$5,000+$8,000=$22,600\begin{array}{c}\\{\rm{Total manufacturing cost}} = \left( {{\rm{Direct material}}} \right) + \left( {{\rm{Direct labor}}} \right) + \left( {{\rm{Manufacturing overhead}}} \right)\\\\ = \$ 9,600 + \$ 5,000 + \$ 8,000\\\\ = \$ 22,600\\\end{array}

(a.2)

Compute the ending work-in-progress (WIP) inventory of case 1 as given below:

Costofgoodsmanufactured=(Totalmanufacturingcost)+(BeginningWIPinventory)(EndingWIPinventory)EndingWIPinventory=(Totalmanufacturingcost)+(BeginningWIPinventory)(Costofgoodsmanufactured)=$22,600+$1,000$17,000=$6,600\begin{array}{c}\\{\rm{Cost of goods manufactured}} = \left( {{\rm{Total manufacturing cost}}} \right) + \left( {{\rm{Beginning WIP inventory}}} \right)\\\\ - \left( {{\rm{Ending WIP inventory}}} \right)\\\\{\rm{Ending WIP inventory}} = \left( {{\rm{Total manufacturing cost}}} \right) + \left( {{\rm{Beginning WIP inventory}}} \right)\\\\ - \left( {{\rm{Cost of goods manufactured}}} \right)\\\\ = \$ 22,600 + \$ 1,000 - \$ 17,000\\\\ = \$ 6,600\\\end{array}

(a.3)

Compute the beginning inventory of finished goods in case 1.

Beginningfinishedgoodsinventory=(Goodsavailableforsale)(Costofgoodsmanufactured)=$20,000$17,000=$3,000\begin{array}{c}\\{\rm{Beginning finished goods inventory}} = \left( {{\rm{Goods available for sale}}} \right)\\\\ - \left( {{\rm{Cost of goods manufactured}}} \right)\\\\ = \$ 20,000 - \$ 17,000\\\\ = \$ 3,000\\\end{array}

(a.4)

Compute the cost of inventory sold in case 1 as given below:

Costofgoodssold=(Goodsavailableforsale)(Endingfinishedgoodsinventory)=$20,000$3,400=$16,600\begin{array}{c}\\{\rm{Cost of goods sold}} = \left( {{\rm{Goods available for sale}}} \right) - \left( {{\rm{Ending finished goods inventory}}} \right)\\\\ = \$ 20,000 - \$ 3,400\\\\ = \$ 16,600\\\end{array}

(a.5)

Compute gross profit of case 1 as given below:

Grossprofit=(Salesrevenue)(Salesdiscount)(Costofgoodssold)=$24,500$2,500$16,600=$5,400\begin{array}{c}\\{\rm{Gross profit}} = \left( {{\rm{Sales revenue}}} \right) - \left( {{\rm{Sales discount}}} \right) - \left( {{\rm{Cost of goods sold}}} \right)\\\\ = \$ 24,500 - \$ 2,500 - \$ 16,600\\\\ = \$ 5,400\\\end{array}

(a.6)

Compute net income of case 1 as given below:

Netincome=GrossprofitOperatingexpenses=$5,400$2,500=$2,900\begin{array}{c}\\{\rm{Net income}} = {\rm{Gross profit}} - {\rm{Operating expenses}}\\\\ = \$ 5,400 - \$ 2,500\\\\ = \$ 2,900\\\end{array}

(a.7)

Compute beginning work-in-progress inventory of case 2 as given below:

Costofgoodsmanufactured=(Totalmanufacturingcost)+(BeginningWIPinventory)(EndingWIPinventory)BeginningWIPinventory=(Costofgoodsmanufactured)(Totalmanufacturingcost)+(EndingWIPinventory)=$22,000$16,000+$3,000=$9,000\begin{array}{c}\\{\rm{Cost of goods manufactured}} = \left( {{\rm{Total manufacturing cost}}} \right) + \left( {{\rm{Beginning WIP inventory}}} \right)\\\\ - \left( {{\rm{Ending WIP inventory}}} \right)\\\\{\rm{Beginning WIP inventory}} = \left( {{\rm{Cost of goods manufactured}}} \right) - \left( {{\rm{Total manufacturing cost}}} \right)\\\\ + \left( {{\rm{Ending WIP inventory}}} \right)\\\\ = \$ 22,000 - \$ 16,000 + \$ 3,000\\\\ = \$ 9,000\\\end{array}

(a.8)

Determine the cost of goods available for sale of case 2 as given below:

Goodsavailableforsale=(Costofgoodsmanufactured)+(Beginningfinishedgoodsinventory)=$22,000+$3,300=$25,300\begin{array}{c}\\{\rm{Goods available for sale}} = \left( {{\rm{Cost of goods manufactured}}} \right)\\\\ + \left( {{\rm{Beginning finished goods inventory}}} \right)\\\\ = \$ 22,000 + \$ 3,300\\\\ = \$ 25,300\\\end{array}

(a.9)

Determine the cost of goods sold in case 2 as given below:

Costofgoodssold=(Goodsavailableforsale)(Endingfinishedgoodsinventory)=$25,300$2,500=$22,800\begin{array}{c}\\{\rm{Cost of goods sold}} = \left( {{\rm{Goods available for sale}}} \right) - \left( {{\rm{Ending finished goods inventory}}} \right)\\\\ = \$ 25,300 - \$ 2,500\\\\ = \$ 22,800\\\end{array}

(a.10)

Compute the amount of sales revenue of case 2 as given below:

Salesrevenue=(Costofgoodssold)+(Salesdiscount)+(Grossprofit)=$22,800+$1,400+$7,000=$31,200\begin{array}{c}\\{\rm{Sales revenue}} = \left( {{\rm{Cost of goods sold}}} \right) + \left( {{\rm{Sales discount}}} \right) + \left( {{\rm{Gross profit}}} \right)\\\\ = \$ 22,800 + \$ 1,400 + \$ 7,000\\\\ = \$ 31,200\\\end{array}

(a.11)

Determine operating expenses of case 2 as given below:

Operatingexpenses=GrossprofitNetincome=$7,000$5,000=$2,000\begin{array}{c}\\{\rm{Operating expenses}} = {\rm{Gross profit}} - {\rm{Net income}}\\\\ = \$ 7,000 - \$ 5,000\\\\ = \$ 2,000\\\end{array}

(b)

Prepare the condensed cost of goods manufactured for case 1 as given below:

Schedule of cost of goods manufactured
for case 1
Particulars
Amount
Raw material used
$9,600
Direct labor
$5,000
Manufacturi

Prepare the condensed cost of goods manufactured for case 1 using spreadsheet formulas as given below:

3
1
Schedule of cost of goods manufactured
2
for case 1
Particulars
Amount
4. Raw material used
9600
5 Direct labor
5000
6 Ma

(c.1)

Prepare the income statement for case 1 as given below:

Income statement
for case 1
Particulars Amount
Sales
$24,500
Less: Sales discount $2,500
Net sales
$22,000
Less: Cost of good

Prepare the income statement for case 1 using spreadsheet formulas as given below:

A
В
Income statement
for case 1
Particulars Amount
Sales
24500
Less: Sales discount 2500
Net sales
=C4-C5
7 Less: Cost of goo

(c.2)

Prepare the section of current assets for case 1 as given below:

Schedule of Current assets
for case 1
Particulars
Amount
Cash
$4,000
Receivables (net)
$15,000
Inventory:
Raw material
$600
W

Prepare the section of current assets for case 1 using spreadsheet formulas as given below:

3
Schedule of Current assets
for case 1
Particulars
Amount
4 Cash
4000
5 Receivables (net)
15000
6 Inventory:
7 Raw material

Ans: Part a.1

The total manufacturing cost of case 1 is $22,600.

Part a.2

Ending work-in-progress inventory of case 1 is $6,600.

Part a.3

Beginning finished goods inventory in case 1 is $3,000.

Part a.4

Cost of goods sold in case 1 is $16,600.

Part a.5

Gross profit for case 1 is $5,400.

Part a.6

Net income of case 1 is $2,900.

Part a.7

Beginning work-in-progress inventory of case 2 is $9,000.

Part a.8

Goods available for sale of case 2 is $25,300.

Part a.9

Cost of goods sold in case 2 is $22,800.

Part a.10

Sales revenue of case 2 is $31,200.

Part a.11

Operating expenses for case 2 are $12,000.

Part b

Cost of goods manufactured for case 1 is $17,000.

Part c.1

Net income of case 1 is $2,900.

Part c.2

Current assets for case 1 is $30,000.